OpenStudy (anonymous):

A child's box of bricks contains cubes, cones, and spheres. Two cones and a sphere on one side of a pair of scale will just balance a cube on the other side; and a sphere and a cube together will just balance three cones. How many spheres will just balance a single cone?

OpenStudy (epoweritheta):

C +S =c S+c=3C adding C+2S+c=c+3C C=S C-wt of Cone c-wt of cube S-wt of sphere